# Application of Whisper in Clinical Practice: the Post-Stroke Speech Assessment during a Naming Task thanks:

### v1

@@ III-E Evaluation metrics
The verbatim transcriptions of the SONIVA-Naming dataset trials often consisted of more than the target Naming task word. A target word detection accuracy was calculated by assessing whether the predicted transcriptions correctly identified the target word exclusively. A prediction was classified as True Positive when the target word was present in both the ground truth transcription and prediction, and True Negative when it was absent in both. The presence of the target word in the ground truth label but not in the prediction indicated a False Negative and the reverse - a False Positive.
III-F Statistical Analysis
@@
The effect of model type (i.e., baseline vs fine-tuned) and size (i.e., small vs medium) on WER across all trials was assessed on the test set using the Friedman test, followed by Mann-Whitney U tests with Bonferroni correction for multiple comparisons. The non-parametric nature of the tests was due to the non-normal distributions of the WER within each dataset.
@@
To complete the analysis, all categorical variables (e.g., Sex, Smoking status, English as second language, Previous Stroke), were one-hot encoded. The distribution of continuous variables (LDL Cholesterol, speech fluency) was tested for normality using the Shapiro-Wilk test. Since the assumption of normality was met, continuous variables between groups were compared with a Student’s t-test. Categorical variables were compared between groups using Fisher’s exact test. Statistical significance was set at $p<0.05$.
IV Results
IV-A Verbatim Transcription
A significant effect of model size and type on WER was detected when testing on synthetic ($\chi^{2}$(11, $N$=42) = 313.95, $p<0.001$), SONIVA-Naming healthy ($\chi^{2}$(11, $N$=806) = 5215.60, $p<0.01$), and SONIVA-Naming patient datasets ($\chi^{2}$(11, $N$=524) = 2442.80, $p<0.01$). All fine-tuned models significantly outperformed baseline Whisper in transcription accuracy on the synthetic, healthy and patient datasets ($p<0.01$) (Table II). Fine-tuning on healthy and patient speech resulted in 87.72% and 71.22% improvement in WER for healthy and patient speech compared to baseline model of the same size. A significant improvement was also observed in case of synthetic data, where the best-performing models were Medium ft-h, Medium ft-p and Small and Medium ft-all, reaching a WER of 0%, compared to a WER of minimum 85.71% in case of Whisper baseline. The best-performing model on the healthy data was the Small ft-h, reaching average WER of 8.82%. On the patient dataset, Medium ft-hp reached the lowest average WER of 26.35%. However, this Medium model was not significantly better than its Small counterpart ($p>0.05$). Additionally, when evaluated on the patient dataset, the difference in performance between ft-hp and single-dataset trained models ft-h and ft-p was insignificant ($p>0.05$) in both model sizes.
On the unseen TORGO database, a significant effect of model size and type on WER was also detected ($\chi^{2}$(11, $N$=1240) = 850.08, $p<0.001$) (Table III). All fine-tuned models significantly outperformed baseline Whisper in verbatim transcription accuracy ($p<0.01$). The best performing model was Medium ft-syn, which achieved a 22.64% reduction in WER compared to the baseline model in the same size. However, it did not perform significantly better than other fine-tuned models ($p>0.5$), with the exception of Medium ft-hp ($p=0.045$).
TABLE II: Comparison of word error rate in transcription across testing datasets of SONIVA-Naming data.
@@ IV-A Verbatim Transcription
Size
Testing Dataset Model Small Medium
Synthetic data Baseline 92.85 85.71
ft-syn 45.23 21.42
ft-h 2.38 0
ft-p 7.14 0
ft-hp 2.38 2.38
ft-all 0 0
SONIVA-Naming healthy Baseline 96.54 90.87
ft-syn 53.64 48.06
ft-h 8.82 9.89
ft-p 14.60 11.94
ft-hp 12.69 12.58
ft-all 12.00 9.07
SONIVA-Naming patient Baseline 100.54 97.57
ft-syn 75.83 71.53
ft-h 28.38 27.72
ft-p 28.82 26.80
ft-hp 27.79 26.35
ft-all 27.04 29.15
TABLE III: Comparison of word error rate in transcription on the out-of-domain TORGO database
Word Error Rate (%)
Size
Testing Dataset Model Small Medium
TORGO database Baseline 97.82 96.13
ft-syn 77.72 73.49
ft-h 77.39 75.65
ft-p 75.01 74.62
ft-hp 78.92 79.67
ft-all 78.10 76.58
IV-B Target Word Detection accuracy
Compared to baseline Whisper, fine-tuned models had a better performance on target word detection accuracy in both healthy and patient speech (Table IV). For healthy speech, the best-performing model was the Small ft-h, reaching an accuracy of 0.97. This increase in performance was due to the ft-h model making noticeably fewer False Negative mistakes ($n=27$) compared to baseline Whisper ($n=496$). On the other hand, for patient speech, the best-performing models were the Small and Medium ft-hp, reaching an accuracy of 0.92. This model made fewer False Negative mistakes ($n=39$) than the Medium baseline model ($n=286$). However, Small and Medium ft-hp models produced 2 False Positive mistakes, while the Small and Medium baseline Whisper only had 1 and 0 False Positive mistakes, respectively. Overall, an improvement of up to 61% and 51% was observed for healthy and patient data, respectively.
TABLE IV: Model target word detection performance on the SONIVA-Naming dataset
Testing Dataset Size Model Accuracy True Positive True Negative False Positive False Negative
@@ IV-B Target Word Detection accuracy
IV-C Accuracy Score Prediction
The best-performing models for predicting the Semantic, Dysfluency, and Self-correction accuracy scores during the Naming task on healthy data were the Medium models initialized with ft-h weights, achieving F1 Macro scores of 0.7449, 0.8390, and 0.7539, respectively (Table V). For the Phonology metric, the best-performing model was the Small model also initialized with ft-h weights, reaching an F1 Macro of 0.6424. Compared to their respective baselines, an average improvement of 7.24 ± 4.01% was observed across all accuracy metrics. Although all models had moderate to high F1 Macro scores, there was a discrepancy between their performance on correct/unimpaired (class 1) and incorrect/impaired (class 0) trials. The F1 scores for class 1 trials were 0.9814, 0.9564, 0.9860, and 0.9744, whereas the F1 scores for class 0 trials were 0.5085, 0.7215, 0.5217, and 0.3103 for the Semantic, Dysfluency, Phonology, and Self-correction metric models, respectively.
In context of patient speech, the best-performing model for predicting Semantic accuracy scores was the Small model initialised with ft-p encoder weights, reaching F1 Macro of 0.7659. For predicting Dysfluency and Self-correction, the best-performing models were the Small models initialized with baseline encoder weights, reaching F1 Macro of 0.9021 and 0.7112, respectively (Table V). For Phonology, the best model was the Medium model initialized with ft-p encoder weights, achieving F1 Macro of 0.6435. The observed improvement of fine-tuned models in Phonology and Semantics was in average 4.425 ± 0.175%. In Dysfluency and Self-correction, the baseline models performed on average 1.45±0.06% better than corresponding models initialized with ft-p encoder weights. While the discrepancy between performance on class 1 and class 0 trials was still evident, it was lower than on healthy data. The F1 scores in class 1 trials were 0.9105, 0.9282, 0.8944 and 0.8604, whereas F1 scores in class 0 were 0.6213, 0.8759, 0.5279 and 0.4265 for Semantic, Dysfluency, Self-correction and Phonology metric models, respectively.
TABLE V: F1 Macro for Different Accuracy Metrics and Testing Datasets
Testing Dataset Size Encoder Weight Semantic Dysfluency Self-correction Phonology
@@ IV-C Accuracy Score Prediction
IV-D Predictive Validity Analysis
@@
Patients identified as impaired based on model predictions had significantly lower speech fluency than those identified as unimpaired based on Semantic ($t(16)=-3.61,p=0.01$), Dysfluency ($t(16)=-2.3880,p=0.02$) and Phonology ($t(16)=-2.3880,p=0.02$) metrics. Significantly more impaired patients spoke English as a second language compared to unimpaired patients, based on Dysfluency and Phonology metrics ($p=0.02$). Additionally, a higher number of patients with previous stroke history were identified as impaired than unimpaired based on Semantic ($p=0.02$), Dysfluency ($p=0.04$) and Phonology ($p=0.04$) metrics. The divergent validity analysis showed no significant differences between impaired and unimpaired patients for age, sex, LDL cholesterol or smoking status, as expected $(p>0.05)$.

### v2

@@ 3.5 Evaluation metrics
Verbatim transcriptions of the SONIVA-Naming dataset trials often contained words beyond the target Naming task word. A target word detection accuracy was calculated by assessing whether the predicted transcriptions correctly identified the target word only. A prediction was classified as a True Positive if the target word appeared in both the ground truth and prediction and a True Negative if absent in both. The presence of the target word in the ground truth label but not in the prediction indicated a False Negative and the reverse - a False Positive.
3.6 Statistical Analysis
@@
The effect of model type (i.e., baseline vs fine-tuned) and size (i.e., Small vs Medium) on WER across all trials was assessed on the test set using the Friedman test, followed by Mann-Whitney U tests with Bonferroni correction for multiple comparisons. The non-parametric tests were used due to the non-normal distributions of the WER within each dataset.
@@
To complete the analysis, all categorical variables (e.g., sex, smoking status, English as second language, previous stroke history), were one-hot encoded. The distribution of continuous variables (LDL cholesterol, speech fluency) was tested for normality using the Shapiro-Wilk test. Since the assumption of normality was met, continuous variables between groups were compared with a Student’s t-test. Categorical variables were compared between groups using Fisher’s exact test.
4 Results
4.1 Verbatim Transcription
A significant effect of model size and type on WER was detected when testing on synthetic ($\chi^{2}$(11, NN=42) = 313.95, $p<0.001$), SONIVA-Naming healthy ($\chi^{2}$(11, NN=806) = 5215.60, $p<0.01$), and SONIVA-Naming patient datasets ($\chi^{2}$(11, NN=524) = 2442.80, $p<0.01$). All fine-tuned models significantly outperformed baseline Whisper in transcription accuracy on the synthetic, healthy and patient datasets ($p<0.01$) (Table 1). Fine-tuning on healthy and patient speech resulted in 87.72% and 71.22% improvement in WER for healthy and patient speech compared to baseline model of the same size. A significant improvement was also observed in case of synthetic data, where the best-performing models were Medium ft-h, Medium ft-p and Small and Medium ft-all, reaching a WER of 0%, compared to a WER of minimum 85.71% in case of Whisper baseline. The best-performing model on the healthy data was the Small ft-h, reaching average WER of 8.82%. On the patient dataset, Medium ft-hp reached the lowest average WER of 26.35%. However, this Medium model was not significantly better than its Small counterpart ($p>0.05$). Additionally, when evaluated on the patient dataset, the difference in performance between ft-hp and single-dataset trained models ft-h and ft-p was insignificant ($p>0.05$) in both model sizes.
On the unseen TORGO database, a significant effect of model size and type on WER was also detected ($\chi^{2}$(11, NN=1240) = 850.08, $p<0.001$) (Table 1). All fine-tuned models significantly outperformed baseline Whisper in verbatim transcription accuracy ($p<0.01$). The best performing model was Medium ft-syn, which achieved a 22.64% reduction in WER compared to the baseline model in the same size. However, it did not perform significantly better than other fine-tuned models ($p>0.5$), with the exception of Medium ft-hp ($p=0.045$).
Table 1: Comparison of word error rate (WER) across testing datasets.
@@ 4.1 Verbatim Transcription
Model  Size  Synthetic
SONIVA
healthy
SONIVA
patient
TORGO
Baseline    Small    92.85    96.54    100.54    97.82
  Medium    85.71    90.87    97.57    96.13
ft-syn    Small    45.23    53.64    75.83    77.72
  Medium    21.42    48.06    71.53    73.49
ft-h    Small    2.38    8.82  28.38    77.39
  Medium    0  9.89    27.72    75.65
ft-p    Small    7.14    14.60    28.82    75.01
  Medium    0  11.94    26.80    74.62
ft-hp    Small    2.38    12.69    27.79    78.92
  Medium    2.38    12.58    26.35  79.67
ft-all    Small    0  12.00    27.04    78.10
  Medium    0  9.07    29.15    76.58
4.2 Target Word Detection accuracy
Compared to baseline Whisper, fine-tuned models had a higher target word detection accuracy in both healthy and patient speech (Table 2). For healthy speech, the best-performing model was the Small ft-h, reaching an accuracy of 0.97. This increase in performance was due to the ft-h model making noticeably fewer False Negative mistakes ($n=27$) compared to baseline Whisper ($n=496$). For patient speech, the best-performing models were the Small and Medium ft-hp, reaching an accuracy of 0.92. This model made fewer False Negative mistakes ($n=39$) than the Medium baseline model ($n=286$). However, Small and Medium ft-hp models produced 2 False Positive mistakes, compared to 1 and 0 False Positive mistakes in Small and Medium baseline models, respectively. Overall, an improvement of up to 61% and 51% was observed for healthy and patient data, respectively.
Table 2: Model target word detection performance on the SONIVA-Naming dataset.
Testing Dataset  Size  Model  Accuracy
True
Positive
True
Negative
False
Positive
False
Negative
@@ 4.2 Target Word Detection accuracy
4.3 Accuracy Score Prediction
The best-performing models for predicting the Semantic, Dysfluency, and Self-correction accuracy scores during the Naming task on healthy data were the Medium models initialized with ft-h weights, achieving F1 Macro scores of 0.7449, 0.8390, and 0.7539, respectively (Table 3). For the Phonology metric, the best-performing model was the Small model also initialized with ft-h weights, reaching an F1 Macro of 0.6424. Compared to their respective baselines, an average improvement of 7.24 ± 4.01% was observed across all accuracy metrics. Although all models had moderate to high F1 Macro scores, there was a discrepancy between their performance on correct/unimpaired (class 1) and incorrect/impaired (class 0) trials. The F1 scores for class 1 trials were 0.9814, 0.9564, 0.9860, and 0.9744, whereas the F1 scores for class 0 trials were 0.5085, 0.7215, 0.5217, and 0.3103 for the Semantic, Dysfluency, Phonology, and Self-correction metric models, respectively.
In context of patient speech, the best-performing model for predicting Semantic accuracy scores was the Small model initialised with ft-p encoder weights, reaching F1 Macro of 0.7659. For predicting Dysfluency and Self-correction, the best-performing models were the Small models initialized with baseline encoder weights, reaching F1 Macro of 0.9021 and 0.7112, respectively (Table 3). For Phonology, the best model was the Medium model initialized with ft-p encoder weights, achieving F1 Macro of 0.6435. The observed improvement of fine-tuned models in Phonology and Semantics was in average 4.425 ± 0.175%. In Dysfluency and Self-correction, the baseline models performed on average 1.45±0.06% better than corresponding models initialized with ft-p encoder weights. While the discrepancy between performance on class 1 and class 0 trials was still evident, it was lower than on healthy data. The F1 scores in class 1 trials were 0.9105, 0.9282, 0.8944 and 0.8604, whereas F1 scores in class 0 were 0.6213, 0.8759, 0.5279 and 0.4265 for Semantic, Dysfluency, Self-correction and Phonology metric models, respectively.
Table 3: F1 Macro for different accuracy metrics and testing datasets.
Testing Dataset  Size  Encoder Weight  Semantic  Dysfluency
Self-
correction
Phonology
@@ 4.3 Accuracy Score Prediction
4.4 Predictive Validity Analysis
@@
Patients identified as impaired based on model predictions had significantly lower speech fluency than those identified as unimpaired based on Semantic (t(16)=3.61,p=0.01t(16)=-3.61,p=0.01), Dysfluency (t(16)=2.3880,p=0.02t(16)=-2.3880,p=0.02) and Phonology (t(16)=2.3880,p=0.02t(16)=-2.3880,p=0.02) metrics. Significantly more impaired patients spoke English as a second language compared to unimpaired patients, based on Dysfluency and Phonology metrics ($p=0.02$). Additionally, a higher number of patients with previous stroke history were identified as impaired than unimpaired based on Semantic ($p=0.02$), Dysfluency ($p=0.04$) and Phonology ($p=0.04$) metrics. The divergent validity analysis showed no significant differences between impaired and unimpaired patients for sex, LDL cholesterol or smoking status, as expected $(p>0.05)$ (See Appendix E).
5 Discussion
The baseline Whisper model performed substantially worse on the verbatim transcription of single-word SONIVA-Naming healthy speech compared to previously reported results on healthy speech. Previous studies have reported WER as low as 3.40% and 2.90% for the baseline Whisper Small and Medium models, respectively [8]. However, these results were obtained using the LibriSpeech corpus [35], a standard ASR benchmark consisting of continuous speech from audiobook recordings. In contrast, the markedly poorer performance observed in the current study likely reflects the short-form nature of the single-word Naming task-derived speech, which provides limited contextual information for decoding and therefore reduces transcription accuracy.
Although the synthetic speech dataset was free from natural speech variability and background noise, the baseline Whisper models also performed poorly on this dataset (WER = 92.85% and 85.71% for the Small and Medium models, respectively). The observed performance further supports that Whisper does not generalize directly to short-form speech and highlights the necessity of task-specific fine-tuning. Further, once fine-tuned, all models achieved their best performance on the synthetic dataset, as anticipated, since it represented the linguistically and acoustically simplest condition.
All fine-tuned Whisper models improved transcription performance across datasets. Models trained on healthy and patient data further improved the WER compared to those trained on synthetic data, suggesting the importance of natural speech diversity in improving transcription performance.
The fine-tuned model showed adequate performance for post-stroke speech transcription. In patients, the best performing model was the Medium ft-hp model achieving a 26.35% WER. However, this result was not significantly different from Medium ft-p and ft-h models (26.80% WER ; $p>0.05$ and 27.72% WER ; $p>0.05$, respectively). Similarly, target detection performance of Medium ft-hp, ft-p and ft-h models was also comparable (0.92, 0.91 and 0.91 accuracy, respectively).
This suggests that fine-tuning on age-matched, task-specific healthy speech may generalize sufficiently to patient data for the same task. This is particularly relevant in clinical applications since collection of healthy speech data is more accessible, and can facilitate a larger dataset for fine-tuning.
Further, comparable performance between ft-h, ft-p and ft-hp models on patient speech suggests that the main performance advantage comes from fine-tuning Whisper on short-form speech produced in the single-word Naming task, irrespective of the data source.
The WER achieved by the ft-p model was consistent with previous finding on wav2vec 2.0 and Whisper models fine-tuned on aphasic speech (WER 22.30%–55.50% depending on aphasia severity and 21.93% WER, respectively) [20, 29], and outperformed hyperparameter-tuned Whisper (WER 38.50%) [26]. However, as prior work targeted continuous speech, direct comparison is limited.
Medium models achieved the best performance across most datasets but did not differ significantly in WER from the smaller variants. Given the higher computational cost and training time, Small models may offer a more efficient option for future clinical use, particularly when scaling to larger datasets.
Similar to the SONIVA-Naming test set, fine-tuned Whisper models showed markedly improved transcription on the unseen TORGO dataset compared to the baseline, underscoring the need to adapt pretrained models for out-of-domain clinical data. Nonetheless, the best-performing model (Medium ft-syn) still produced a high WER of 73.49%, well above clinical usability. Models fine-tuned on SONIVA patient data also performed poorly on TORGO (Medium WER: 74.62%). This can be attributed to clinical heterogeneity: SONIVA targets stroke-induced aphasia involving both motoric and linguistic processing difficulties, whereas TORGO focuses on dysarthria and motoric impairments only. These results emphasize the importance of tailoring pretrained models to the specific linguistic and pathological features of each condition to achieve robust generalization and clinical utility.
In terms of target detection accuracy on the SONIVA-Naming task data, fine-tuned models showed higher True Positive and lower False Negative scores compared to baseline Whisper, improving clinical applicability. However, on patient data, the fine-tuned models produced more False Positive mistakes, as a result of failing to transcribe phonological errors made by patients (e.g., predicting target word ‘comb’ instead of ‘comg’). As Whisper is designed to predict the next most probable token in the sequence, these discrepancies are expected.
Whisper was also used to complete a stroke-specific downstream task, aimed at estimating the level of speech impairment. In case of healthy data, the models initialized with ft-h encoder weights had the best F1 Macro in accuracy score prediction, suggesting the importance of fine-tuning. However, for patient data, models initialized with ft-p encoder weights improved performance on the Semantic and Phonology metrics but not on Dysfluency and Self-correction. Specifically, while the fine-tuned model outperformed previous studies that used Whisper encoder features for classifying dysfluencies in stuttering (F1 Macro of 0.71) [23], its performance was slightly lower than the baseline for Dysfluency. This reduction in performance suggests that while fine-tuning for verbatim transcription captures linguistic features like semantics and phonology, it may not capture more complex patterns in patient speech, such as hesitations and self-corrections, which are less vital for transcription but crucial for Dysfluency and Self-correction predictions, leading to worse performance on these metrics for patient speech. It is, however, worth noting that the absolute difference between the best baseline encoder models and their fine-tuned encoder versions was negligible (1.45±0.06%).
Despite this difference between model performance on impaired and unimpaired trials, the patient -level clinical validity analysis confirmed that the predicted accuracy scores were clinically meaningful. A significant difference between patients identified as impaired or unimpaired based on model predictions was observed in speech fluency, English proficiency status, and previous stroke history, as expected [37, 38, 39]. The results of the divergent validity analysis showed no differences in features unrelated to speech impairment such as cholesterol levels, sex and smoking status.
6 Limitations and Future Work
We note several limitations and directions for future work. First, patients were not stratified by speech impairment severity when splitting into training, validation, and test sets, which may have led to uneven distributions and limited generalizability across impairment levels. Nonetheless, this should not affect the main conclusions about Whisper, as similar findings were observed in healthy data.
In the accuracy prediction task, performance could be improved by addressing class imbalance through multiple binary classification tasks (instead of a multi-class one) and by using approaches alternative to downsampling. For example, data perturbation, modifying audio features such as pitch, frequency, or tempo, to create more samples has been used in disordered and aphasic speech transcription [42, 43, 41]. Alternatively, in-domain data augmentation, such as incorporating similar tasks from datasets like AphasiaBank [11], could also be applied [40].
Whisper’s transcription accuracy could also be evaluated using beam search instead of greedy decoding, to create a more complete assessment of its applicability to stroke-specific data. Greedy decoding selects the most probable token at each step in the output sequence, whereas beam search considers multiple candidate sequences and chooses the final output based on cumulative probability [8, 44].
In order to obtain a broader overview on the application of foundation models to pathological speech in stroke, Whisper should be compared with other transformer-based architectures, such as Seamless M4T v2, a multimodal model capable of speech transcription that has shown competitive results compared to Whisper on ASR tasks [45]. Finally, future work should explore whether token-level loss analyses, encoder probing, or modality-specific pretraining can help disentangle the extent to which current models capture generalizable acoustic markers of clinical speech, particularly in settings where contextual compensation from the decoder is limited, such as single-word utterances.
7 Conclusion
We evaluated the applicability of Whisper, a general foundation model, to a specific clinical case study on a single-word speech task in patients with stroke. Our findings confirmed that additional fine-tuning on single-word data is required to achieve efficient transcription performance that can facilitate clinically useful downstream tasks. Further, comparisons of fine-tuned model performance on in-domain and out-of-domain neurological disorder datasets suggests further need for model adaptation to disorder-specific data. With such fine-tuning we can advance the automated analysis of pathological speech allowing for more efficient diagnosis and monitoring of stroke.
Appendix A Sample demographics of the SONIVA-Naming dataset
Table 4: Sample demographics of the SONIVA-Naming dataset.
Demographic Feature  SONIVA-Naming healthy  SONIVA-Naming patients
  N = 132    N = 87
  Mean (Standard Deviation)
Age    61.6 (10.8)    61.8 (13.6)
Sex∗∗∗
    Male:Female    52:77    62:25
    Missing data    3    0
English Language
    Non-native    27    30
    Native    102    57
Education level∗∗∗
    School    36    45
    Degree    43    37
    Post-graduate    37    3
    Missing data    16    2
$^{*}p<0.05$;  $^{***}p<0.001$
Appendix B Whisper Architecture
To complete the accuracy classification task, WhisperforAudioClassification model configuration [32] was used, which retains the same Whisper encoder architecture, but replaces the decoder with a linear classifier projection head, consisting of two linear layers.
Appendix C The general study pipeline
Figure 1: General study pipeline. Overview of the training and inference pipeline for verbatim transcription and accuracy prediction models. Transcription: Training – Whisper models were fine-tuned on speech data from 5 different datasets. A synthetic dataset of audio-transcription word pairs based on stimuli in the Naming task, SONIVA-Naming healthy dataset, collected from healthy age-matched participants performing the Naming task, SONIVA-Naming patient dataset, collected from patients with stroke performing the Naming task, a combined dataset of SONIVA-Naming healthy and patient data, and a combined dataset of all available data (synthetic, SONIVA-Naming healthy and SONIVA-Naming patient). All models were trained using cross-entropy loss to transcribe spoken words (e.g. “acorn”) from the Naming task input audio. Inference – Fine-tuned models (ft-syn, ft-h, ft-p, ft-hp and ft-all) were evaluated on 4 separate datasets – synthetic, SONIVA-Naming healthy, SONIVA-Naming patient, and an unseen testing dataset of dysarthric speech derived from the TORGO database. Accuracy: Training – accuracy prediction models were trained on SONIVA-Naming derived healthy and patient speech in a linear probing framework. The encoder layer of the models was frozen and the classification projection head was trained to predict accuracy scores using cross-entropy loss (e.g “2"). The encoder weights of the models were initialized at baseline or with the weights derived from the models trained for verbatim transcription of healthy or patient speech
Appendix D Additional training details
All model training was conducted on a single NVIDIA RTX 6000 GPU. Mixed-precision arithmetic (fp16) and gradient checkpointing was implemented to improve computational efficiency and optimise memory usage. Model training and evaluation were conducted using the PyTorch framework [46] and HuggingFace Transformers library [47].
Appendix E Predictive Validity Analysis